OpenStudy (anonymous):

find the slope, if it exists, of the line containing the pair of points. (5,0) and (0,-5) the slope m= or the slope is undefined?

OpenStudy (anonymous):

\[m={y_{1}-y_{2} \over x_{1}-x_{2}}\]

OpenStudy (anonymous):

I'd use the "rise over run" traditional slope equation:\[m=\frac{y_2-y_1}{x_2-x_1}\]You are given two points to use, so label one of those points (x1,y1) and the other as (x2,y2) (order doesn't mater), and plug in and solve!

OpenStudy (anonymous):

someone answered and said slope undefined Since you are given an X= coordinate only, it goes straight up and down. If you were given Y=coordinate, the slope would be zero.

myininaya (myininaya):

(-5-0)/(0-5)=-5/(-5)=1 slope is 1

OpenStudy (anonymous):

Not true...you are given both x and y for both points. Point #1 sits on the x-axis to the right, and point #2 sits on the y-axis downward. By definition they can't be vertical to each other, so there IS a slope...

myininaya (myininaya):

if you have (5,2) and (6,2) *see that the y's are the same in bot so slope will be zero if you have (3,4) and (3,9) * see that the x's are the same in both so slope is undefined

OpenStudy (anonymous):

so my question is is the answer undefined or is the slope 1

OpenStudy (anonymous):

The slope is 1. Only vertical lines have undefined slope. Vertical lines have identical x-values with different y-values. In this example, that's not the case.

OpenStudy (anonymous):

londovi & myininaya the problem reads only: x=-5 find the slope if it exists or undefined

myininaya (myininaya):

it ask find the slope of the line containing two points: (5,0) and (0,-5)

myininaya (myininaya):

equation for this line is y=x-5

myininaya (myininaya):

not x=5 and not x=-5

OpenStudy (anonymous):

If the problem reads as x = -5, then you have yourself a vertical line (all vertical lines have the form of x = constant), and that would mean you're back to an undefined slope.

OpenStudy (anonymous):

myininaya & londovir the problem states x = -5 I am looking at my paper and that is how it reads there is no y involved

OpenStudy (anonymous):

thanks londovir that is exactly how it reads you get the medal

myininaya (myininaya):

lol you should make a new thread if you are asking a different question from this person who posted this question i thought we have been talking about the question that this thread states

OpenStudy (anonymous):

@jazi1: Thanks and you're welcome. @myininaya: I know, I was a little confused too! :-)

OpenStudy (anonymous):

i got way confused...lol

myininaya (myininaya):

your slope is 1 lish k not undefined lol

OpenStudy (anonymous):

thanks all of you for trying

OpenStudy (anonymous):

ah, thank you everyone!